OpenStudy (anonymous):

which expression is the first step you take when finding the quotient of 5/9 divided by 3/10? 9/5 * 3/10 9/5 * 10/3 5/9 * 10/3

OpenStudy (anonymous):

9/5 * 10/3

OpenStudy (anonymous):

but how'd you get that?~

OpenStudy (anonymous):

i divided

OpenStudy (lgbasallote):

@lalalalala993 take the reciprocal of the second fraction

OpenStudy (lgbasallote):

then multiply

OpenStudy (lgbasallote):

for example \[\frac 12 \div \frac 23 \implies \frac 12 \times \frac 32\]

OpenStudy (lgbasallote):

do you get it?

OpenStudy (anonymous):

do u get it

OpenStudy (anonymous):

@lalalalala993

OpenStudy (anonymous):

co it would be 5/9 * 10/3 then, right? @lgbasallote

OpenStudy (anonymous):

*so

OpenStudy (lgbasallote):

absolutely

OpenStudy (anonymous):

thanks for the help!

OpenStudy (anonymous):

welcome